OpenStudy (anonymous):

I'm really not sure whether or not I solved this problem correctly! h(x)=x^2-9 f(x)=2x+8 So what is h(f(x))? Would it be h(f(x))= 1(2x+8)?

OpenStudy (mathmate):

One way to look at it is: h(x)=x^2-9 y=f(x)=2x+8 h(f(x))=h(y)=(y^2)-9=...

OpenStudy (anonymous):

Thank you for the help!

OpenStudy (mathmate):

You're welcome! Feel free to post your answer for checking if you wish.